First answer is incorrect, confused on how to solve this.. thanks for the help in advance!
Calculate the present value of the following annuity streams: a. $7,000 received each year for 4 years on the last day of each year if your investments pay 6 percent compounded annually b. $7,00O received each quarter for 4 years on the last day of each quarter if your investments pay 6 percent compounded quarterly c. $7000 received each year for 4 years on the first day of each year if your investments pay 6 percent compounded annually d. $7,00O received each quarter for 4 years on the first day of each quarter itf your investments pay 6 percent compounded quarterly (For all requirements, do not round intermediate calculations. Round your answers to 2 decimal places. (e.g.. 32.16)) 29.486.55 a. Present value b. Present value c. Present value d. Present value
Answer #1

Answer:

a)

Annuity A=7000/ year

r=6%

n=4 years

Present value (PV) =A*(1-(1+r)^-n)/r

PV=7000*(1-(1+6%)^-4)/6%

PV=$24255.74

B)

Annuity A=7000/ quarter

interest rate=6% per year

r=6%/4=1.5% per quarter

years=4

n=4*4=16 quarters

Present value (PV) =A*(1-(1+r)^-n)/r

PV=7000*(1-(1+1.5%)^-16)/1.5%

PV=$98918.85

C)

Annuity A=7000/ year

r=6%

n=4 years

Present value (PV) =A+ A*(1-(1+r)^-(n-1))/r

PV=7000+7000*(1-(1+6%)^-3)/6%

PV=$25711.08

D)

Annuity A=7000/ quarter

interest rate=6% per year

r=6%/4=1.5% per quarter

years=4

n=4*4=16 quarters

Present value (PV) =A+A*(1-(1+r)^-(n-1))/r

PV=7000+7000*(1-(1+1.5%)^-15)/1.5%

PV=$100402.63